# DitHub: A Modular Framework for Incremental Open-Vocabulary Object Detection

### v2

@@ 6 Conclusions
This Appendix provides supplementary information regarding the main manuscript.
  • App. A — Further Insights on Efficiency: Optimizes DitHub’s performance when adapting Grounding DINO.
    • Sec. A.1 — Blockwise Adaptation Analysis: Justifies adapting only the encoder for DitHub based on empirical metrics.
    • Sec. A.2 — Employing Full LoRA for Class-Specificity: Explores class-specific LoRA, favoring a parameter-efficient shared $B$ matrix.
  • App. B — Further Insights on Forgetting: Analyzes DitHub’s anti-forgetting capabilities.
    • Sec. B.1 — Forgetting on ODinW-13: Presents anti-forgetting results for ODinW-13.
    • Sec. B.2 — Forgetting on ODinW-O: Demonstrates robust anti-forgetting on the ODinW-O benchmark.
  • App. C — Additional Benchmark Details: Provides comprehensive dataset information.
    • Sec. C.1 — ODinW-13 and MS COCO: Describes ODinW-13 for specialization and MS COCO (ZCOCO) for zero-shot evaluation.
    • Sec. C.2 — ODinW-O: Introduces the ODinW-O benchmark for Incremental Learning and domain shifts.
  • App. D — Additional Results: Presents complementary experimental information.
    • Sec. D.1 — Standard Deviations: Reports standard deviations for reproduced methods from Tab. 1 and Tab. 2.
    • Sec. D.2 — Unlearning Additionals: Illustrates DitHub’s modular capability for selective knowledge removal (Fig. C).
  • App. E — Implementation Details: Outlines experimental setup, including hardware, model configurations, and optimization specifics.
  • App. F — Limitations and Broader Impact: Discusses current limitations and broader societal implications.
  • App. G — Qualitative Results: Provides visual comparisons (Fig. D) demonstrating DitHub’s improved detection and robustness.
@@ A.1 Blockwise Adaptation Analysis
While Grounding DINO consists of four main components – an image backbone for visual encoding, a textual backbone for language understanding, and an encoder-decoder pair for detection – we focus our adaptation on the latter two. This choice is motivated by the need for a more detection-centric adaptation and efficiency considerations, as the textual backbone, being a BERT Base model [5], accounts for more than half of Grounding DINO ’s parameters. In Tab. A, we present results for adapting only the encoder (Fig. A), only the decoder (Fig. A), and both encoder and decoder (Fig. A). DitHub employs the first approach – adapting only the encoder – as it delivers the best performance on both ZCOCO and Avg metrics.
@@ A.2 Employing Full LoRA for Class-Specificity
Table B: LoRA full results on ODinW-13, with DitHub results taken from Tab. 1 for reference.
@@
While our methodology leverages a shared and unique $B$ matrix for all tasks, Tab. B presents results obtained using class-specific pairs. In this variant of DitHub, both the $A$ and $B$ low-rank matrices of LoRA undergo the same specialized optimization process: an initial warmup phase followed by dedicated specialization. In other words, the training pipeline for matrix $B$ mirrors that of matrix $A$ described in Sec. 4.1. Applying our full set of strategies to both matrices yields a $\boldsymbol{\texttt{+}1.03}$ mAP improvement in average performance across the tasks of ODinW-13, albeit at the cost of doubling the number of trainable parameters (Fig. 3). Prioritizing efficiency, we opt for a lightweight implementation that trades a marginal performance gain for significantly reduced memory requirements. Notably, ZCOCO results favor our efficient approach, as a single, unspecialized $B$ matrix remains more aligned with the pretraining distribution, thereby mitigating degradation in zero-shot capabilities.
@@ Appendix B Further Insights on Forgetting
This section presents an additional analysis of the anti-forgetting properties of DitHub, complementing the study in Sec. 5.2 by offering an alternative perspective. We report the forgetting metric for both ODinW-13 and ODinW-O, computed for DitHub and its primary competitor, ZiRa [4]. This metric is calculated using the canonical definition [3, 1]: the difference in performance on a specific task after learning the entire sequence of tasks compared to the original performance on that task immediately after its initial learning.
@@ B.1 Forgetting on ODinW-13
Table C: Forgetting metrics in terms of mAP for the ODinW-13 dataset. The numbers above each task indicate the order in which they were processed. Bold indicates best results.
$11$ $10$ $2$ $13$ $12$ $1$ $7$ $6$ $3$ $4$ $9$ $8$ $5$
@@
ZiRa -$\texttt{-}6.66$$\texttt{-}4.31$$\texttt{-}4.82$$\boldsymbol{\texttt{-}1.80}$ $0.00$$\texttt{-}2.09$$\texttt{-}8.71$$\texttt{-}5.80$$\texttt{-}7.76$$\texttt{-}12.63$$\texttt{-}19.19$$\boldsymbol{\texttt{-}1.82}$$\texttt{-}9.85$$\texttt{-}7.77$
DitHub -$\boldsymbol{\texttt{-}1.73}$$\boldsymbol{\texttt{-}1.56}$$\boldsymbol{\texttt{-}1.69}$$\texttt{-}2.34$ $0.00$$\boldsymbol{\texttt{-}1.61}$ $\boldsymbol{~{}~{}3.62}$$\boldsymbol{\texttt{-}1.91}$ $\boldsymbol{~{}~{}1.54}$$\boldsymbol{\texttt{-}6.94}$$\boldsymbol{\texttt{-}7.27}$$\texttt{-}2.11$ $\boldsymbol{~{}~{}0.10}$$\boldsymbol{\texttt{-}2.34}$
Tab. C presents the forgetting in terms of mAP points for the $13$ tasks of ODinW-13, with the average across all tasks reported as Avg. DitHub demonstrates strong performance on both average and individual tasks, consistently outperforming its competitor by a notable margin. Given the limited class overlap within ODinW-13, the primary component susceptible to catastrophic forgetting in this analysis is the shared matrix $B$.
@@ B.2 Forgetting on ODinW-O
Table D: Forgetting metrics in terms of mAP for the ODinW-O dataset. Bold indicates best results.
$5$ $6$ $1$ $2$ $4$ $3$
@@
ZiRa -$\texttt{-}3.84$$\boldsymbol{\texttt{-}0.08}$ $0.00$$\texttt{-}9.67$$\texttt{-}10.66$$\texttt{-}0.98$$\texttt{-}1.62$
DitHub -$\boldsymbol{\texttt{-}2.68}$$\texttt{-}4.75$ $0.00$$\boldsymbol{\texttt{-}2.74}$$\boldsymbol{\texttt{-}9.51}$ $\boldsymbol{~{}~{}1.74}$$\boldsymbol{\texttt{-}0.79}$
Tab. D reports the forgetting metric results for the ODinW-O benchmark. Unlike the previous analysis, this assessment allows us to evaluate the forgetting of DitHub’s class-specific $A$ components for all classes within the dataset. Indeed, ODinW-O is specifically designed to test Incremental Learning capabilities, featuring complete class overlap across tasks. This overlap leads to catastrophic forgetting of the $A_{c}$ expert component for the same class $c$ across different domains. It’s worth noting that even in this setting, which emphasizes the $A$ component, the forgetting of $B$ remains a concern. Even in this challenging context, DitHub surpasses its competitor by an appreciable margin, demonstrating difficulty only with the Ae task and highlighting the non-trivial nature of merging and preserving knowledge from the aerial domain.
@@ C.2 ODinW-O
We introduce a new benchmark, ODinW-O (Overlapped), designed to further evaluate the class-incremental and specialization capabilities of DitHub. Derived from ODinW-35 [19], ODinW-O retains only classes shared across at least one task (i.e., overlapping classes), while removing tasks with no shared classes. This filtering is crucial because retaining classes that do not reappear in subsequent tasks would cause the benchmark to resemble the original one. Specifically, the corresponding non-recurring class modules would remain frozen, requiring no further adaptation. In contrast, ODinW-O eliminates this scenario by ensuring all retained classes appear in multiple tasks. By enforcing strictly overlapping classes, ODinW-O provides a challenging benchmark for evaluating Incremental Learning. Additionally, to rigorously evaluate the model robustness to domain shifts, we designed ODinW-O by explicitly selecting datasets that feature markedly different contexts (Fig. B), enabling clear insights into how effectively the model adapts to variations in visual appearance, environmental conditions, and imaging domains. We select the following datasets, already mentioned in Sec. C.1, and extracted only the overlapping classes:
@@ D.1 Standard Deviations
We report the standard deviations of the reproduced methods from Tab. 1 and Tab. 2 in Tab. E and F, respectively, computed over three runs with different random seeds.
D.2 Unlearning Additionals
As previously discussed in Sec. 5.2, our modular framework facilitates the precise removal of specific knowledge from Grounding DINO through the subtraction of corresponding experts. Fig. C visually represents this process: each column depicts the mean Average Precision (mAP) of a target class after unlearning different modules (reported on the y-axis), with colors normalized column-wise to emphasize relative performance decrements. The prominent darker diagonal pattern signifies that removing a class-specific module results in the most substantial performance reduction for its associated class. This observation substantiates that our modules effectively encapsulate specialized knowledge, as their unlearning selectively impacts their respective classes. Furthermore, this matrix representation elucidates correlations between certain classes during the unlearning procedure. For instance, the darker shading in the “bicycle” column indicates that the "person" class exhibits the second most significant mAP drop. This correlation arises from the frequent co-occurrence of “person” instances riding bicycles, naturally linking these classes within the comprehensive unlearning pipeline.
Table E: Standard deviations of the reproduced methods in Tab. 1.
@@ D.2 Unlearning Additionals
1 ZiRa ±$\pm 0.90$$\pm 2.05$$\pm 1.49$$\pm 0.17$$\pm 2.16$$\pm 0.98$$\pm 0.79$$\pm 1.63$$\pm 1.33$$\pm 7.91$$\pm 3.29$$\pm 12.17$$\pm 4.45$$\pm 3.07$$\pm 3.25$
DitHub ±$\pm 0.60$$\pm 0.81$$\pm 0.95$$\pm 0.57$$\pm 2.66$$\pm 1.58$$\pm 3.72$$\pm 0.00$$\pm 1.53$$\pm 1.10$$\pm 2.90$$\pm 7.03$$\pm 3.68$$\pm 4.66$$\pm 2.46$
5 ZiRa ±$\pm 0.23$$\pm 0.43$$\pm 3.76$$\pm 2.15$$\pm 1.40$$\pm 1.08$$\pm 3.54$$\pm 3.74$$\pm 1.54$$\pm 4.26$$\pm 3.73$$\pm 4.13$$\pm 2.03$$\pm 2.02$$\pm 4.16$
DitHub ±$\pm 0.42$$\pm 1.27$$\pm 2.52$$\pm 0.70$$\pm 2.42$$\pm 6.74$$\pm 0.76$$\pm 3.27$$\pm 0.17$$\pm 5.49$$\pm 4.79$$\pm 2.22$$\pm 4.29$$\pm 1.81$$\pm 1.54$
10 ZiRa ±$\pm 0.28$$\pm 0.66$$\pm 0.44$$\pm 2.64$$\pm 2.77$$\pm 0.52$$\pm 6.43$$\pm 0.97$$\pm 1.04$$\pm 2.17$$\pm 2.04$$\pm 3.90$$\pm 1.21$$\pm 3.71$$\pm 0.87$
DitHub ±$\pm 0.20$$\pm 0.54$$\pm 2.26$$\pm 0.72$$\pm 4.23$$\pm 3.53$$\pm 1.35$$\pm 3.19$$\pm 0.30$$\pm 2.74$$\pm 0.34$$\pm 2.99$$\pm 1.01$$\pm 1.15$$\pm 1.27$
Full ZiRA ±$\pm 0.44$$\pm 0.68$$\pm 2.73$$\pm 1.17$$\pm 1.39$$\pm 4.77$$\pm 7.26$$\pm 2.19$$\pm 1.33$$\pm 0.33$$\pm 4.14$$\pm 3.57$$\pm 2.44$$\pm 2.82$$\pm 3.13$
DitHub ±$\pm 0.16$$\pm 0.44$$\pm 2.39$$\pm 1.00$$\pm 2.83$$\pm 0.48$$\pm 1.34$$\pm 0.60$$\pm 0.38$$\pm 1.68$$\pm 0.76$$\pm 1.57$$\pm 2.87$$\pm 0.69$$\pm 1.01$
Table F: Standard deviations of the reproduced methods in Tab. 2.
@@
ZiRa ±$\pm 0.25$$\pm 0.27$$\pm 1.76$$\pm 1.14$$\pm 1.35$$\pm 0.32$$\pm 1.75$$\pm 1.76$
DitHub ±$\pm 0.24$$\pm 0.66$$\pm 2.39$$\pm 1.88$$\pm 0.76$$\pm 1.31$$\pm 0.46$$\pm 0.84$

### v3

@@ 6 Conclusions
Acknowledgments and Disclosure of Funding
Chiara Cappellino’s research was supported by the STORE project (https://edf-store.com), funded by the European Defence Fund (EDF) under grant agreement EDF-2022-101121405-STORE. Angelo Porrello was financially supported by the Italian Ministry for University and Research – through the ECOSISTER ECS 00000033 CUP E93C22001100001 project – and the European Commission under the Next Generation EU programme PNRR. Rita Cucchiara was financially supported by the EU Horizon project “ELIAS - European Lighthouse of AI for Sustainability” (No. 101120237).
@@
This appendix provides supplementary information to the main manuscript, organized as follows.
Appendix A — Further Insights on Efficiency.
Analyzes the performance optimization of DitHub when adapting Grounding DINO.
  • Section˜A.1 — Blockwise Adaptation Analysis: Justifies adapting only the encoder for DitHub based on empirical metrics.
  • Section˜A.2 — Employing Full LoRA for Class-Specificity: Explores class-specific LoRA, ultimately favoring a parameter-efficient shared $B$ matrix.
  • Section˜A.3 — Performance-Memory Tradeo-ff Details: Details the per-task mAP for the experiments introduced in the right panel of Figure˜7.
Appendix B — Further Insights on Forgetting.
Evaluates the anti-forgetting capabilities of DitHub.
  • Section˜B.1 — Forgetting on ODinW-13: Presents anti-forgetting results for the ODinW-13 benchmark.
  • Section˜B.2 — Forgetting on ODinW-O: Demonstrates robust anti-forgetting performance on the ODinW-O benchmark.
Appendix C — Additional Benchmark Details.
Provides comprehensive information about the datasets used.
  • Section˜C.1 — ODinW-13 and MS COCO: Describes the ODinW-13 dataset for specialization and MS COCO (ZCOCO) for zero-shot evaluation.
  • Section˜C.2 — ODinW-O: Introduces the ODinW-O benchmark for Incremental Learning and domain shift analysis.
Appendix D — Additional Results.
Presents complementary experimental results.
  • Section˜D.1 — Standard Deviations: Reports standard deviations for the reproduced methods in Table˜1 and Table˜2.
  • Section˜D.2 — Further Unlearning Results: Illustrates the modular capability of DitHub for selective knowledge removal, as shown in Figure˜C.
  • Section˜D.3 — Performance with Swin-B Backbone: Shows the scaling capabilities of DitHub when employing a larger backbone.
  • Section˜D.4 — Beyond ODinW-13: Presents results for the AODRaw benchmark.
Appendix E — Implementation Details.
Outlines the experimental setup, including hardware, model configurations, and optimization specifics.
Appendix F — Limitations and Broader Impact.
Discusses the current limitations of the proposed method and its broader societal implications.
Appendix G — Qualitative Results.
Provides visual comparisons in Figure˜D that demonstrate the improved detection accuracy and robustness of DitHub.
@@ A.1 Blockwise Adaptation Analysis
While Grounding DINO consists of four main components – an image backbone for visual encoding, a textual backbone for language understanding, and an encoder-decoder pair for detection – we focus our adaptation on the latter two. This choice is motivated by the need for a more detection-centric adaptation and efficiency considerations, as the textual backbone, being a BERT Base model [5], accounts for more than half of Grounding DINO ’s parameters. In Table˜A, we present results for adapting only the encoder (Figure˜A), only the decoder (Figure˜A), and both encoder and decoder (Figure˜A). DitHub employs the first approach – adapting only the encoder – as it delivers the best performance on both ZCOCO and Avg metrics.
@@ A.2 Employing Full LoRA for Class-Specificity
Table B: LoRA full results on ODinW-13, with DitHub results taken from Table˜1 for reference.
@@
While our methodology leverages a shared and unique $B$ matrix for all tasks, Table˜B presents results obtained using class-specific pairs. In this variant of DitHub, both the $A$ and $B$ low-rank matrices of LoRA undergo the same specialized optimization process: an initial warmup phase followed by dedicated specialization. In other words, the training pipeline for matrix $B$ mirrors that of matrix $A$ described in Section˜4.1. Applying our full set of strategies to both matrices yields a $\boldsymbol{\texttt{+}1.03}$ mAP improvement in average performance across the tasks of ODinW-13, albeit at the cost of doubling the number of trainable parameters (Figure˜3). Prioritizing efficiency, we opt for a lightweight implementation that trades a marginal performance gain for significantly reduced memory requirements. Notably, ZCOCO results favor our efficient approach, as a single, unspecialized $B$ matrix remains more aligned with the pretraining distribution, thereby mitigating degradation in zero-shot capabilities.
A.3 Performance-Memory Trade-off Details
Table C: Performance of DitHub at different ranks on the ODinW-13 benchmark.
Method ZCOCO Avg Ae Aq Co Eg Mu Pa Pv Pi Po Ra Sh Th Ve
DitHub $r=1$ $47.14$ $57.04$ $31.89$ $37.01$ $67.93$ $72.58$ $43.63$ $53.76$ $69.80$ $67.86$ $42.92$ $72.49$ $40.31$ $76.11$ $65.27$
DitHub $r=2$ $46.66$ $60.25$ $35.86$ $49.59$ $71.93$ $74.27$ $54.37$ $63.37$ $67.61$ $70.50$ $39.65$ $64.05$ $49.39$ $76.32$ $66.41$
DitHub $r=4$ $46.73$ $60.53$ $38.54$ $48.59$ $71.13$ $74.45$ $50.46$ $61.49$ $68.20$ $69.97$ $41.48$ $64.97$ $55.64$ $78.13$ $63.84$
DitHub $r=8$ $46.81$ $61.93$ $37.38$ $50.19$ $71.02$ $74.37$ $53.34$ $66.09$ $68.83$ $70.16$ $49.31$ $66.31$ $53.55$ $78.64$ $65.85$
DitHub $r=16$ $47.01$ $62.19$ $34.62$ $50.65$ $70.46$ $68.56$ $49.28$ $65.57$ $69.58$ $71.10$ $56.65$ $70.88$ $52.82$ $79.30$ $68.18$
Table˜C details the per-task mAP for the experiments introduced in Figure˜7 (right).
@@ Appendix B Further Insights on Forgetting
This section presents an additional analysis of the anti-forgetting properties of DitHub, complementing the study in Section˜5.2 by offering an alternative perspective. We report the forgetting metric for both ODinW-13 and ODinW-O, computed for DitHub and its primary competitor, ZiRa [4]. This metric is calculated using the canonical definition [3, 1]: the difference in performance on a specific task after learning the entire sequence of tasks compared to the original performance on that task immediately after its initial learning.
@@ B.1 Forgetting on ODinW-13
Table D: Forgetting metrics in terms of mAP for the ODinW-13 dataset. The numbers above each task indicate the order in which they were processed. Bold indicates best results.
$11$ $10$ 22 $13$ $12$ 11 77 66 33 44 99 88 55
@@
ZiRa -$\texttt{-}6.66$ -$\texttt{-}4.31$ -$\texttt{-}4.82$ -$\boldsymbol{\texttt{-}1.80}$ $0.00$ -$\texttt{-}2.09$ -$\texttt{-}8.71$ -$\texttt{-}5.80$ -$\texttt{-}7.76$ -$\texttt{-}12.63$ -$\texttt{-}19.19$ -$\boldsymbol{\texttt{-}1.82}$ -$\texttt{-}9.85$ -$\texttt{-}7.77$
DitHub -$\boldsymbol{\texttt{-}1.73}$ -$\boldsymbol{\texttt{-}1.56}$ -$\boldsymbol{\texttt{-}1.69}$ -$\texttt{-}2.34$ $0.00$ -$\boldsymbol{\texttt{-}1.61}$ $\boldsymbol{~~3.62}$ -$\boldsymbol{\texttt{-}1.91}$ $\boldsymbol{~~1.54}$ -$\boldsymbol{\texttt{-}6.94}$ -$\boldsymbol{\texttt{-}7.27}$ -$\texttt{-}2.11$ $\boldsymbol{~~0.10}$ -$\boldsymbol{\texttt{-}2.34}$
Table˜D presents the forgetting in terms of mAP points for the $13$ tasks of ODinW-13, with the average across all tasks reported as Avg. DitHub demonstrates strong performance on both average and individual tasks, consistently outperforming its competitor by a notable margin. Given the limited class overlap within ODinW-13, the primary component susceptible to catastrophic forgetting in this analysis is the shared matrix $B$.
@@ B.2 Forgetting on ODinW-O
Table E: Forgetting metrics in terms of mAP for the ODinW-O dataset. Bold indicates best results.
55 66 11 22 44 33
@@
ZiRa -$\texttt{-}3.84$ -$\boldsymbol{\texttt{-}0.08}$ $0.00$ -$\texttt{-}9.67$ -$\texttt{-}10.66$ -$\texttt{-}0.98$ -$\texttt{-}1.62$
DitHub -$\boldsymbol{\texttt{-}2.68}$ -$\texttt{-}4.75$ $0.00$ -$\boldsymbol{\texttt{-}2.74}$ -$\boldsymbol{\texttt{-}9.51}$ $\boldsymbol{~~1.74}$ -$\boldsymbol{\texttt{-}0.79}$
Table˜E reports the forgetting metric results for the ODinW-O benchmark. Unlike the previous analysis, this assessment allows us to evaluate the forgetting of DitHub’s class-specific $A$ components for all classes within the dataset. Indeed, ODinW-O is specifically designed to test Incremental Learning capabilities, featuring complete class overlap across tasks. This overlap leads to catastrophic forgetting of the $A_{c}$ expert component for the same class $c$ across different domains. It’s worth noting that even in this setting, which emphasizes the $A$ component, the forgetting of $B$ remains a concern. Even in this challenging context, DitHub surpasses its competitor by an appreciable margin, demonstrating difficulty only with the Ae task and highlighting the non-trivial nature of merging and preserving knowledge from the aerial domain.
@@ C.2 ODinW-O
We introduce a new benchmark, ODinW-O (Overlapped), designed to further evaluate the class-incremental and specialization capabilities of DitHub. Derived from ODinW-35 [19], ODinW-O retains only classes shared across at least one task (i.e., overlapping classes), while removing tasks with no shared classes. This filtering is crucial because retaining classes that do not reappear in subsequent tasks would cause the benchmark to resemble the original one. Specifically, the corresponding non-recurring class modules would remain frozen, requiring no further adaptation. In contrast, ODinW-O eliminates this scenario by ensuring all retained classes appear in multiple tasks. By enforcing strictly overlapping classes, ODinW-O provides a challenging benchmark for evaluating Incremental Learning. Additionally, to rigorously evaluate the model robustness to domain shifts, we designed ODinW-O by explicitly selecting datasets that feature markedly different contexts (Figure˜B), enabling clear insights into how effectively the model adapts to variations in visual appearance, environmental conditions, and imaging domains. We select the following datasets, already mentioned in Section˜C.1, and extracted only the overlapping classes:
@@ D.1 Standard Deviations
We report the standard deviations of the reproduced methods from Table˜1 and Table˜2 in Table˜F and G, respectively, computed over three runs with different random seeds.
D.2 Further Unlearning Results
As previously discussed in Section˜5.2, our modular framework facilitates the precise removal of specific knowledge from Grounding DINO through the subtraction of corresponding experts. Figure˜C visually represents this process: each column depicts the mean Average Precision (mAP) of a target class after unlearning different modules (reported on the y-axis), with colors normalized column-wise to emphasize relative performance decrements. The prominent darker diagonal pattern signifies that removing a class-specific module results in the most substantial performance reduction for its associated class. This observation substantiates that our modules effectively encapsulate specialized knowledge, as their unlearning selectively impacts their respective classes. Furthermore, this matrix representation elucidates correlations between certain classes during the unlearning procedure. For instance, the darker shading in the “bicycle” column indicates that the "person" class exhibits the second most significant mAP drop. This correlation arises from the frequent co-occurrence of “person” instances riding bicycles, naturally linking these classes within the comprehensive unlearning pipeline.
Table F: Standard deviations of the reproduced methods in Table˜1.
@@ D.2 Further Unlearning Results
1 ZiRa $\pm 0.90$ $\pm 2.05$ $\pm 1.49$ $\pm 0.17$ $\pm 2.16$ $\pm 0.98$ $\pm 0.79$ $\pm 1.63$ $\pm 1.33$ $\pm 7.91$ $\pm 3.29$ $\pm 12.17$ $\pm 4.45$ $\pm 3.07$ $\pm 3.25$
DitHub $\pm 0.60$ $\pm 0.81$ $\pm 0.95$ $\pm 0.57$ $\pm 2.66$ $\pm 1.58$ $\pm 3.72$ $\pm 0.00$ $\pm 1.53$ $\pm 1.10$ $\pm 2.90$ $\pm 7.03$ $\pm 3.68$ $\pm 4.66$ $\pm 2.46$
5 ZiRa $\pm 0.23$ $\pm 0.43$ $\pm 3.76$ $\pm 2.15$ $\pm 1.40$ $\pm 1.08$ $\pm 3.54$ $\pm 3.74$ $\pm 1.54$ $\pm 4.26$ $\pm 3.73$ $\pm 4.13$ $\pm 2.03$ $\pm 2.02$ $\pm 4.16$
DitHub $\pm 0.42$ $\pm 1.27$ $\pm 2.52$ $\pm 0.70$ $\pm 2.42$ $\pm 6.74$ $\pm 0.76$ $\pm 3.27$ $\pm 0.17$ $\pm 5.49$ $\pm 4.79$ $\pm 2.22$ $\pm 4.29$ $\pm 1.81$ $\pm 1.54$
10 ZiRa $\pm 0.28$ $\pm 0.66$ $\pm 0.44$ $\pm 2.64$ $\pm 2.77$ $\pm 0.52$ $\pm 6.43$ $\pm 0.97$ $\pm 1.04$ $\pm 2.17$ $\pm 2.04$ $\pm 3.90$ $\pm 1.21$ $\pm 3.71$ $\pm 0.87$
DitHub $\pm 0.20$ $\pm 0.54$ $\pm 2.26$ $\pm 0.72$ $\pm 4.23$ $\pm 3.53$ $\pm 1.35$ $\pm 3.19$ $\pm 0.30$ $\pm 2.74$ $\pm 0.34$ $\pm 2.99$ $\pm 1.01$ $\pm 1.15$ $\pm 1.27$
Full ZiRA $\pm 0.44$ $\pm 0.68$ $\pm 2.73$ $\pm 1.17$ $\pm 1.39$ $\pm 4.77$ $\pm 7.26$ $\pm 2.19$ $\pm 1.33$ $\pm 0.33$ $\pm 4.14$ $\pm 3.57$ $\pm 2.44$ $\pm 2.82$ $\pm 3.13$
DitHub $\pm 0.16$ $\pm 0.44$ $\pm 2.39$ $\pm 1.00$ $\pm 2.83$ $\pm 0.48$ $\pm 1.34$ $\pm 0.60$ $\pm 0.38$ $\pm 1.68$ $\pm 0.76$ $\pm 1.57$ $\pm 2.87$ $\pm 0.69$ $\pm 1.01$
Table G: Standard deviations of the reproduced methods in Table˜2.
@@
ZiRa $\pm 0.25$ $\pm 0.27$ $\pm 1.76$ $\pm 1.14$ $\pm 1.35$ $\pm 0.32$ $\pm 1.75$ $\pm 1.76$
DitHub $\pm 0.24$ $\pm 0.66$ $\pm 2.39$ $\pm 1.88$ $\pm 0.76$ $\pm 1.31$ $\pm 0.46$ $\pm 0.84$
D.3 Performance with Swin-B Backbone
Table H: Performance Evaluation with the Swin-B backbone on ODinW-13.
Method ZCOCO Avg Ae Aq Co Eg Mu Pa Pv Pi Po Ra Sh Th Ve
GD ZS w/Swin-B $56.47$ $59.82$ $33.32$ $43.26$ $72.98$ $74.90$ $37.67$ $59.14$ $70.30$ $73.56$ $48.42$ $67.30$ $62.61$ $70.69$ $64.48$
ZiRa w/Swin-B $53.22$ $65.09$ $40.71$ $54.16$ $73.32$ $73.97$ $47.25$ $67.03$ $72.39$ $73.29$ $59.83$ $73.05$ $61.98$ $82.59$ $66.63$
DitHub w/Swin-B $53.82$ $67.03$ $42.41$ $53.74$ $74.53$ $75.18$ $53.28$ $66.05$ $74.04$ $74.22$ $59.25$ $76.51$ $67.63$ $82.85$ $71.67$
To further assess the scalability of our approach, we conducted new experiments utilizing the larger, publicly available pre-trained Grounding DINO model with a Swin-B backbone. Table˜H details the comparative performance of our method, DitHub, against the primary competitor, ZiRa. The findings confirm that DitHub scales effectively, maintaining its robust performance when integrated with a more powerful backbone model.
D.4 Beyond ODinW-13
Table I: Domain-wise performance comparison on ODinW-13.
Method ZCOCO Avg fog_low_light fog_rain fog indoor_low_light indoor_normal_light low_light_rain low_light normal_light rain
GD ZS $47.40$ $15.65$ $13.32$ $19.56$ $19.42$ $15.84$ $24.77$ $11.74$ $16.28$ $5.81$ $14.11$
ZiRA $44.39$ $32.34$ $27.60$ $36.34$ $29.28$ $28.28$ $43.18$ $30.67$ $31.73$ $30.49$ $33.47$
DitHub $45.23$ $34.03$ $30.30$ $35.48$ $31.40$ $31.02$ $47.77$ $31.36$ $32.88$ $30.90$ $35.12$
To rigorously evaluate the generalizability of our approach, we have extended our experiments to a more diverse benchmark. Our initial experimental setup necessitated a benchmark with class re-occurrence across semantically distinct domains — a characteristic not commonly found in standard long-tailed datasets.
For this purpose, we selected AODRaw [22], a challenging and recently introduced benchmark. This dataset is ideally suited for our evaluation, as it comprises $9$ distinct domains with substantial class overlap, encompassing a wide range of meteorological conditions as well as both indoor and outdoor scenes. The performance analysis is presented in Table˜I.